OpenStudy (anonymous):

The harvest increased by 120 percent. If the harvest is 125 bushels, what was the harvest last year ?

OpenStudy (anonymous):

120 % is 1.20 in decimal form Now last year harvest increased by 120 percent equals 125 bushels Let x represent last years bushel so 1.20x=125 Solve for x

OpenStudy (amistre64):

now = p + p(.20) seems reasonable to me

OpenStudy (anonymous):

amistre that would give u now=p(1+.20) now=p(1.20) now=1.20p

OpenStudy (anonymous):

i got it thanks!

OpenStudy (anonymous):

great :D

OpenStudy (anonymous):

can you help me with another?

OpenStudy (anonymous):

sure